# Dynamics of Streamers and Pseudostreamers and Implications for the Solar Wind

### v1

@@ 2.3 Energetics
To sustain the million Kelvin order temperature of the solar corona, there are several possible mechanisms to consider in numerical models, e.g., Alfvénic turbulence, random heating, and damping of magnetohydrodynamic waves. In our setup, we choose a simple volumetric heating function ($H$) that has been used in a number of prior studies (Fan2017; Gannouni2023; Maity2024):
@@
Here $H_{0}={2.86\times 10^{-6}~\rm erg\cdot cm^{-3}\cdot s^{-1}}$, $\lambda_{0}=0.7$ $R_{\odot}$ Mm, $H_{1}={5.72\times 10^{-5}~\rm erg\cdot cm^{-3}\cdot s^{-1}}$, $\lambda_{1}=40$ Mm and $r_{0}=1R_{\odot}$. The first term of the heating function contributes to maintaining the coronal temperature on a global scale (Gannouni2023) by introducing total power of $8.44\times 10^{27}$ $\rm erg\cdot s^{-1}$. The second source term generates the relatively hot closed magnetic fields compared to the open field regions (lionello2009s), which are frequently detected in coronal observations. The total injected power for this comparatively short length scale heating source is $1.37\times 10^{28}$ $\rm erg\cdot s^{-1}$.
It is well established that maintaining realistic energy fluxes across the solar transition region is hugely challenging computationally, due to the abrupt change in temperature from the chromosphere to the corona. Under-resolving the transition region leads to an underestimation of the conductive energy flux, with the result that the energy balance does not match the true coronal conditions. To mitigate these issues, we employ the Transition Region Adaptive Conduction method (TRAC; jhonston2019; jhonston2020) to artificially broaden the transition region while maintaining the appropriate energy balances. Specifically, we use the TRAC implementation from  jhonston2020 in MPI-AMRVAC.
3 Plasma Characteristics in quasi-equilibrium state
@@
Using the simulation setup and physics described above, we evolve the system from the initial state until it reaches a quasi-steady state at 80 hours of solar time. Until $t=20$ hours, we restrict the maximum refinement to five levels while multiple large-scale shock waves pass through in response to the strong initial force-imbalanced state. As shown in Figures 1 and 2(a), the quasi-equilibrium state exhibits helmet streamer and pseudostreamer structures at Northern and Southern latitudes, respectively. The temperature distribution of the whole domain at t=80 hours is shown in Figure 1, which has many localized structures in contrast to the initial uniform 1.25 MK temperature. It is clearly noticeable that the plasma near the magnetic poles is cooler than the closed magnetic field regions of both streamers. At $r=2$ R, the temperature at the north pole is 1.39 MK, where the closed field regions of the helmet streamer and pseudostreamer are at 2.16 MK and 1.65 MK, respectively. In Figure. 3, the temperature profiles further support this contrast between the polar region and the Helmet streamer till $r=10~R_{\odot}$ distance. It is important to mention that the rate of energy-density injection as the coronal heating process (discussed in Section 2.3) is the same for all latitudes at a given radius. Therefore, this temperature contrast between the open and closed magnetic field regions is mostly due to the magnetic field-directed thermal heat flux. In the open magnetic field regions near the magnetic poles, heat can easily escape, and hence, the local plasma becomes cool. For the closed magnetic fields, heat flux is trapped between magnetic footpoints and temperature remains relatively higher than in open field regions. Similarly, the electron density profile ($n$) in the open and closed field regions also exhibits a strong variation as shown in Figure. 2(b): for example, at $3R_{\odot}$, Helmet streamer density is 7.31 times denser than the polar region. This same contrast approximately holds even at higher distances (e.g., $r=10~R_{\odot}$) when measured above the streamer stalk.
@@
Figure 3: Variation of the radial flow speed and temperature of the solar wind with height (at t=80 hours). Black and blue curves represent radial outflow speed (left axis) measured near the north pole and along the helmet streamer stalk. respectively. Temperature profile (right axis) near the polar region and along the same streamer stalk are shown with green and red lines, respectively.
@@
In terms of the plasma kinematics, solar wind streams are developed from the initial zero-velocity state and sustained self-consistently. In our model, coronal gas pressure drives the outflow as the wind. For a Sun-like G-type star, it is believed to be the primary source of wind generation (Lamers1999). In observations, the solar wind speed varies significantly with solar latitude (e.g., mccomas2008). We observe a similar qualitative behaviour of the wind speed profile in our model. At $15$ R, the synthetic wind obtains 261 km/s radial speed near the polar regions (see Figure 3). At a similar height, the wind above the Helmet streamer stalk reaches approximately 245 km/s. In a distance range between $7R_{\odot}$ and $10~R_{\odot}$, plasma from the streamer stalk outflows faster than the polar counterpart. This is because of a propagating local eruption from the helmet streamer, which we will discuss in subsection 4.1. The solar wind is further accelerated and reaches a speed of 320 km/s at $30$ R. Inside the closed field structures of both streamers, plasma motion is dictated by the magnetic field geometry; hence, there are no signatures of the wind. However, continuously fluctuating field-aligned flows are present, with speeds up to 8 km/s in multiple magnetic loops of both streamers. We speculate that these field-aligned flows, together with the thermodynamic evolution, are responsible for the dynamic evolution, particularly of the pseudosteamer, discussed in the following section.
@@
After reaching the quasi-steady state, we introduce a shearing velocity profile near one of the legs of each of the streamers on the lower boundary. The aim of the shearing velocity perturbation in the azimuthal direction ($v_{\phi}$) is to introduce a similar effect as the supergranulation motion in the dynamics, introducing additional free energy into the magnetic field. The driving shearing profile is constant in time and applied for approximately four supergranulation life cycles (96 hours). The explicit form of the perturbation for the helmet streamer ($v^{h}_{\phi}$) is defined in Equ. (12)
@@
where $v_{0}$, $\theta_{1}$ and $\delta\theta_{1}$ represent the amplitude of the shear, location of the helmet streamer foot point in the quasi-steady state at t=80 hours) and width of the shear profile, respectively. We set $v_{0}=0.5$ km/s, $\theta_{1}=44.1^{\circ},\,\delta\theta_{1}=1.71^{\circ}$. Similarly, for the pseudostreamer, the shear velocity ($v^{p}_{\phi}$) is described in Equ. (13)
@@
4 Dynamics of the Helmet Streamer and Pseudostreamer
@@
Figure 4: Quasi-periodic dynamics of the Helmet streamer: The temperature map from the quasi-equilibrium state of the Helmet streamer is shown in panel (a). An arrow is located at $r=16$ R to show the location for synthetic in-situ measurements, discussed in Section 5. Panels (b)–(d) represent different phases of streamer dynamics after introducing the shear in the azimuthal $(\phi)$ direction.
@@
Figure 5: Dynamics in the vicinity of the pseudostreamer, revealing the presence of Interchange Magnetic Reconnection (IMR). The radial component of the magnetic field is represented by the colour map and the magnetic field lines are traced from random seed points. Similar to Fig 4, the quasi-equilibrium state is shown in panel (a) and panels (b)–(d) refer to different times during the time period in which the shear is applied. An arrow is located at $r=2.5$ R for a reference to compute synthetic in-situ observables (see Section 5).
4.1 Helmet streamer dynamics
@@
The evolution of the helmet streamer is clearly noticeable in both the quasi-equilibrium and in the driving phase. A snapshot of the streamer structure from the quasi-steady state is presented in Figure. 4a, while (b)–(d) panels show the different times during the period when the shear driving is applied. The apex of the closed field lines of the helmet streamer in the quasi-steady state is at around $5~R_{\odot}$. However, this height is constantly fluctuating, typically between $5~R_{\odot}-6~R_{\odot}$, while sometimes reaching even larger radii. It is in part modulated by periodic magnetic reconnection events in the overlying Helmet streamer stalk (i.e., in the heliospheric current sheet or HCS). This process leads to the formation of outgoing plasmoid structures as well as sunward reconnection flows or at least decrements in the local outflow speed. After each of these pinching-off reconnection events, the closed magnetic loops of the streamer shrink down to a lower height. With time, gas pressure again builds up in the closed field region due to the heating process and the streamer expands. Subsequently, the solar wind acts to stretch out and eventually open up these closed field lines. The net effect is a “breathing” cycle of the Helmet streamer in which the closed flux contracts and expands, and plasmoids are released along the HCS. The period of this “pinching off” of plasmoids in our simulations is between 11.6414.5511.64-14.55 hours, and leads to a similar periodicity in density fluctuations in the vicinity of the HCS.
To understand the effect of the numerical resolution on the result, we have run an additional simulation in which we restrict the grid refinement to only 5 levels. We find a similar qualitative behaviour in terms of the plasmoid eruptions along the HCS. In the simulation with a coarser grid, the periodicity of the eruption increases compared to the standard simulation (with seven levels of refinement): likely, the reconnection onset is earlier due to the lower effective Lundquist number.
@@
In terms of the plasmoid dynamics following their formation, these structures propagate radially outward with the background solar wind while at the same time expanding. The combined effect of the gas and magnetic pressure inside the plasmoids drives them to expand, while the total pressure in the open field falls off with increasing radius. The plasmoids are magnetically isolated island structures; hence, conductive heat flux cannot equalise their temperature with their surroundings. As a result, they undergo adiabatic cooling while expanding. The temperature can drop from 1.56 MK to 0.66 MK in between $6.8~R_{\odot}-24~R_{\odot}$. We found that for slightly different choices of the coronal heating profile (such as applying Equation (11) without the $r^{-2}$ dependency), the stripping of flux from the helmet streamer can be even more dramatic (see also Suess96). We return to discuss this aspect further in Section 6.
4.2 Pseudostreamer dynamics
@@
Turning now to the evolution of the Pseudostreamer, one of the most prominent features is the presence of Interchange Magnetic Reconnection (IMR) near the base of the streamer stalk, at $r\approx 1.75~R_{\odot}$ – see Figure 5. The two closed magnetic arcades that form the base of the Pseudostreamer are in general, asymmetric in size, as the magnetic flux is not equal at the footpoints of the streamer. In general, the closed loops from one half the streamer are longer than the other and therefore involve higher coronal heating. As a result, the longer closed field lines become hotter, expand, and are compressed against the oppositely oriented open magnetic field. This forms a current sheet, evident from the reversal of the radial component of $\bf B$} across this layer. As a result, IMR takes place. In Figure. 5(a), a small-scale plasmoid is apparent at the upper end of the current sheet, confirming the presence of IMR in this magnetic configuration. During this interchange reconnection, open and closed magnetic field lines reconnect with one another without any net opening or closing of flux. After IMR occurs, part of the previously closed magnetic flux becomes open and exactly the same amount of open flux becomes closed. This distinguishes the reconnection taking place at the pseudostreamer from that at the Helmet streamer stalk. As the IMR removes flux from one closed arcade and adds it to the other, the latter starts to expand, choking off and eventually reversing the orientation of the current sheet, inducing IMR on the opposite flank of the pseudostreamer, and the whole process repeats quasi-periodically. In Figure. 5 (b)–(d) panels, different phases of the IMR can be observed. We observe this repeated reconnection process both during the quasi-steady and driving phase, although the properties are different. In the absence of the shear driving, IMR is solely induced by the interplay of local coronal heating, conduction and radiative cooling process. In the driving phase, both local thermodynamics and shear driving contribute to exciting IMR, and the reconnection process becomes more dynamic, with the formation of a greater number of plasmoids. By driving the shear flow, local magnetic fields become stressed and magnetic free energy builds up. A part of this free energy is dissipated by the IMR process and in the process, plasmoids are ejected. Although there are quasi-periodic plasmoid ejections from the reconnection site, they cannot be advected by the background solar wind for a large distance. The background open magnetic field of the solar wind interacts with the plasmoid’s field, reconnects with one side of the plasmoid and they are quickly eroded away and disappear. It is noteworthy that for exciting and sustaining IMR through the Tearing mode instability, a necessary condition is to reach beyond the critical Lundquist number, which is of the order of $10^{4}$ (biskamp1986; Loureiro2007; Bhattacharjee2009; Gannouni2023). In our model, we find the presence of IMR when the refinement level is increased from 5 to 7 levels, i.e., when the local grid spacing is reduced from 1 Mm to 0.25 Mm.
5 Synthetic In-Situ observation at streamers

### v2

@@ II.3 Energetics
To sustain the million Kelvin order temperature of the solar corona, there are several possible mechanisms to consider in numerical models, e.g., Alfvénic turbulence, random heating, and damping of magnetohydrodynamic waves. In our setup, we choose a simple volumetric heating function ($H$) that has been used in a number of prior studies (Fan, 2017; Gannouni et al., 2023; Sankar Maity et al., 2024; Singh et al., 2025):
@@
Here $H_{0}={2.86\times 10^{-6}~\rm erg\cdot cm^{-3}\cdot s^{-1}}$, $\lambda_{0}=0.7$ $R_{\odot}$ Mm, $H_{1}={5.72\times 10^{-5}~\rm erg\cdot cm^{-3}\cdot s^{-1}}$, $\lambda_{1}=40$ Mm and $r_{0}=1R_{\odot}$. The first term of the heating function contributes to maintaining the coronal temperature on a global scale (Gannouni et al., 2023) by introducing total power of $8.44\times 10^{27}$ $\rm erg\cdot s^{-1}$. The second source term generates the relatively hot closed magnetic fields compared to the open field regions (Lionello et al., 2009), which are frequently detected in coronal observations. The total injected power for this comparatively short length scale heating source is $1.37\times 10^{28}$ $\rm erg\cdot s^{-1}$.
It is well established that maintaining realistic energy fluxes across the solar transition region is hugely challenging computationally, due to the abrupt change in temperature from the chromosphere to the corona. Under-resolving the transition region leads to an underestimation of the conductive energy flux, with the result that the energy balance does not match the true coronal conditions. To mitigate these issues, we employ the Transition Region Adaptive Conduction method (TRAC; Johnston and Bradshaw, 2019; Johnston et al., 2020) to artificially broaden the transition region while maintaining the appropriate energy balances. Specifically, we use the TRAC implementation from  Johnston et al. (2020) in MPI-AMRVAC.
III Plasma Properties of the Quasi-equilibrium State
@@
Using the simulation setup and physics described above, we evolve the system from the initial state until it reaches a quasi-steady state at 80 hours of solar time. Until $t=20$ hours, we restrict the maximum refinement to five levels while multiple large-scale shock waves pass through in response to the strong initial force-imbalanced state. As shown in Figures 1 and 2(a), the quasi-equilibrium state exhibits helmet streamer and pseudostreamer structures at Northern and Southern latitudes, respectively. The temperature distribution of the whole domain at t=80 hours is shown in Figure 1, which has many localized structures in contrast to the initial uniform 1.25 MK temperature (in the coronal part of the domain). It is clearly noticeable that the plasma near the magnetic poles is cooler than the closed magnetic field regions of both streamers. At $r=2$ R, the temperature at the north pole is 1.39 MK, where the closed field regions of the helmet streamer and pseudostreamer are at 2.16 MK and 1.65 MK, respectively. In Figure. 3, the temperature profiles further support this contrast between the polar region and the Helmet streamer out to $r=10~R_{\odot}$.
It is important to mention that the rate of energy-density injection as the coronal heating process (discussed in Section II.3) is the same for all latitudes at a given radius. Therefore, this temperature contrast between the open and closed magnetic field regions is due to the well-known difference in the primary cooling process between closed and open flux regions. In the closed field regions the heating into the corona must be balanced by radiative losses. Thermal conduction serves only to redistribute the energy so that a substantial fraction of the radiative losses are from the transition region (e.g. Johnston and Bradshaw, 2019). As a result, the temperature and especially the density in the closed corona must be high enough to supply the needed losses. In the open field, however, the heating can be balanced by the large enthalpy flux of the solar wind. As a result, the radiative losses are substantially lower, implying lower temperatures and densities. This strong density contrast between the open and closed can be seen in Figure. 2(b): for example, at $3R_{\odot}$, the typical helmet streamer density is 7.31 times larger than in the polar region. This same contrast approximately holds even at higher distances (e.g., $r=10~R_{\odot}$) when measured above the streamer stalk. Note that these results are completely in line with the typical observations.
@@
Figure 3: Variation of the radial flow speed and temperature of the solar wind with height (at t=80 hours). Black and blue curves represent radial outflow speed (left axis) measured near the north pole and along the helmet streamer stalk, respectively. Temperature profile (right axis) near the polar region and along the same streamer stalk are shown with green and red lines, respectively.
@@
In terms of the plasma kinematics, solar wind streams are developed from the initial zero-velocity state and sustained self-consistently. In our model, coronal gas pressure drives the outflow as the wind. For a Sun-like G-type star, it is believed to be the primary source of wind generation (Lamers and Cassinelli, 1999). In observations, the solar wind speed varies significantly with solar latitude (e.g., McComas et al., 2008). We observe a similar qualitative behavior of the wind speed profile in our model. At $15$ R, our wind reaches 261 km/s radial speed near the polar regions (see Figure 3). At a similar height, the wind above the Helmet streamer stalk reaches approximately 245 km/s. In a distance range between $7R_{\odot}$ and $10~R_{\odot}$, plasma from the streamer stalk outflows faster than the polar counterpart. This is because of a propagating local eruption from the helmet streamer, which we will discuss in subsection IV.1. The solar wind is further accelerated and reaches a speed of 320 km/s at $30$ R. Inside the closed field structures of both streamers, plasma motion is dictated by the magnetic field geometry; hence, there are no signatures of the wind. However, continuously fluctuating field-aligned flows are present, with speeds up to 8 km/s in multiple magnetic loops of both streamers. We argue that these field-aligned flows, together with the plasma thermodynamics, are responsible for the dynamic evolution, particularly of the pseudosteamer, discussed in the following section.
@@
After reaching the quasi-steady state, we introduce a shearing velocity profile near one of the legs of each of the streamers on the lower boundary. The aim of the shearing velocity perturbation in the azimuthal direction ($v_{\phi}$) is to simulate the effects on helmet and pseudostreamer and dynamics of the supergranulation motions, which add free energy into the coronal magnetic field. The driving shearing profile is constant in time and applied for approximately four supergranulation life cycles (96 hours). The explicit form of the perturbation for the helmet streamer ($v^{h}_{\phi}$) is defined in Equ. (12)
@@
where $v_{0}$, $\theta_{1}$ and $\delta\theta_{1}$ represent the amplitude of the shear, location of the helmet streamer foot point in the quasi-steady state at t=80 hours, and width of the shear profile, respectively. We set $v_{0}=0.5$ km/s, $\theta_{1}=44.1^{\circ},\,\delta\theta_{1}=1.71^{\circ}$. Similarly, for the pseudostreamer, the shear velocity ($v^{p}_{\phi}$) is described in Equ. (13)
@@
IV Dynamics of the Helmet Streamer and Pseudostreamer
@@
Figure 4: Quasi-periodic dynamics of the Helmet streamer: The temperature map from the quasi-equilibrium state of the Helmet streamer is shown in panel (a). An arrow is located at $r=16$ R to show the location for synthetic in-situ measurements, discussed in Section V. Panels (b)–(d) represent different phases of streamer dynamics after introducing the shear in the azimuthal $(\phi)$ direction.
@@
Figure 5: Dynamics in the vicinity of the pseudostreamer, revealing the presence of Interchange Magnetic Reconnection (IMR). The radial component of the magnetic field is represented by the colour map and the magnetic field lines are traced from random seed points. Similar to Fig 4, the quasi-equilibrium state is shown in panel (a) and panels (b)–(d) refer to different times during the time period in which the shear is applied. An arrow is located at $r=2.5$ R for a reference to compute synthetic in-situ observables (see Section V).
IV.1 Helmet streamer dynamics
@@
The intrinsic non-steady nature of the helmet streamer is clearly evident during both the quasi-equilibrium and driving phases. A snapshot of the streamer structure from the quasi-steady state is presented in Figure. 4a, while (b)–(d) panels show different times during the period when the shear driving is applied. The apex of the closed field lines of the helmet streamer in the quasi-steady state lies around $5~R_{\odot}$. However, this height is constantly fluctuating, typically between $5~R_{\odot}-6~R_{\odot}$, while sometimes reaching even larger radii. It is in part modulated by periodic magnetic reconnection events in the overlying Helmet streamer stalk (i.e., in the heliospheric current sheet or HCS). This process leads to the formation of outgoing plasmoid structures as well as sunward reconnection flows or at least decrements in the local outflow speed. After each of these pinching-off reconnection events, the closed magnetic loops of the streamer shrink down to a lower height. With time, gas pressure again builds up in the closed field region due to the heating process and the streamer expands. Subsequently, the solar wind acts to stretch out and eventually open up these closed field lines. The net effect is a “breathing” cycle of the Helmet streamer in which the closed flux contracts and expands, and plasmoids are released along the HCS. The period of this “pinching off” of plasmoids in our simulations is between 11.6414.5511.64-14.55 hours, and leads to a similar periodicity in density fluctuations in the vicinity of the HCS. This type of continuous plasmoid formation has been seen in other 2.5D (Allred and MacNeice, 2015) and 3D (Higginson and Lynch, 2018) simulations of the quasi-steady HCS, even without the photospheric driving. Furthermore, recent PSP observations very near the Sun, $\sim 10R_{\odot}$, suggest that the HCS is constantly reconnecting and generating plasmoids (Eastwood et al., 2026).
@@
To understand the effect of numerical resolution on the result, we have run an additional simulation in which we restrict the grid refinement to only 5 levels. We find a similar qualitative behaviour in terms of the plasmoid eruptions along the HCS. In the simulation with a coarser grid, the periodicity of the eruption increases compared to the standard simulation (with seven levels of refinement): likely, the reconnection onset is earlier due to the lower effective Lundquist number. In terms of the plasmoid dynamics following their formation, these structures propagate radially outward with the background solar wind while at the same time expanding. Sophisticated observations from the WISPR instrument of the PSP clearly demonstrate the expansion of bright ellipsoidal streamer blob structures (plasmoid in 3D) as they accelerate outward (Wu et al., 2026). In the model, the combined effect of the gas and magnetic pressure inside the plasmoids drives them to expand, while the total pressure in the open field falls off with increasing radius. The plasmoids are magnetically isolated island structures; hence, conductive heat flux cannot equalise their temperature with their surroundings. As a result, they undergo adiabatic cooling while expanding. The temperature can drop from 1.56 MK to 0.66 MK between $6.8~R_{\odot}-24~R_{\odot}$. We found that for slightly different choices of the coronal heating profile (such as applying Equation (11) without the $1/r^{2}$ dependency), the stripping of flux from the helmet streamer can be even more dramatic (see also Suess et al., 1996). We return to discuss this aspect further in Section VI.
IV.2 Pseudostreamer dynamics
@@
Turning now to the evolution of the Pseudostreamer, the most prominent feature is the presence of Interchange Magnetic Reconnection (IMR) near the apex of the pseudostreamer, at $r\approx 1.75~R_{\odot}$ – see Figure 5. The two closed magnetic arcades that form the base of the Pseudostreamer are in general asymmetric in size, as the radial magnetic flux is not symmetric across the photospheric polarity inversion line at the base of the pseudostreamer. In general, the closed loops from one half the streamer will be longer than the other and therefore receive larger total coronal heating. As is well-known from the so-called loop scaling laws, the coronal gas pressure increases with increasing loop length for fixed coronal heating rate (e.g. Rosner et al., 1978). The imbalance in gas pressure causes the closed loops near the pseudostreamer apex, where the plasma beta is high, to expand and compress against the oppositely oriented open magnetic field. This forms a current sheet, evident from the reversal of the radial component of $\bf B$} across this layer. As a result, IMR takes place. In Figure 5(a), a small-scale plasmoid is apparent at the upper end of the current sheet, confirming the presence of IMR in this magnetic configuration. During this interchange reconnection, open and closed magnetic field lines reconnect with one another without any net opening or closing of flux. After IMR occurs, part of the previously closed magnetic flux becomes open and exactly the same amount of open flux becomes closed. This distinguishes the reconnection taking place at the pseudostreamer from that at the Helmet streamer stalk, which involves primarily open-open or closed-closed reconnection, at least for this 2.5D model. As the IMR removes flux from one closed arcade and adds it to the other, the latter starts to expand, choking off and eventually reversing the orientation of the current sheet, inducing IMR on the opposite flank of the pseudostreamer, and the whole process repeats quasi-periodically. In Figure. 5 (b)–(d) panels, different phases of the IMR can be observed. We observe this repeated reconnection process both during the quasi-steady and driving phase, although the properties are different. In the absence of the shear driving, IMR is solely induced by the interplay of local coronal heating, conduction and radiative cooling. In the driving phase, both local thermodynamics and shear driving contribute to exciting IMR, and the reconnection process becomes more dynamic, with the formation of a greater number of plasmoids. By driving the shear flow at the photosphere, magnetic free energy in the form of an out-of-plane ($\phi$ component) field builds up and propagates to the null region where it leads to current sheet formation. A part of this free energy is dissipated by the IMR process and in the process, plasmoids are ejected. It is important to note that even though the driving motions are constant in one direction, the reconnection alternates back and forth between the two closed flux systems. Since the driving speed is very slow compared to the coronal Alfvén speed, then in principle it should be possible for the system to achieve a steady state where the IMR is all in one direction and smooth. Instead, we find bursty reconnection that seems to overshoot, so that the IMR alternates back and forth.
Another key point is that, although there are quasi-periodic plasmoid ejections from the reconnection site, the plasmoids cannot be advected by the background solar wind for a large distance. The background open magnetic field of the solar wind interacts with the plasmoid’s field, reconnects with one side of the plasmoid and they are quickly eroded away and disappear, as observed in many jet models (e.g. Pariat et al., 2009). The net result is to produce a large flux of nonlinear Alfvén waves that carry the energy into the heliosphere, and may well be the origin of the celebrated “Switchbacks” that have been frequently observed by PSP (Wyper et al., 2022). We discuss the properties of these waves further in the following section.
It is noteworthy that for exciting and sustaining reconnection through the tearing mode instability, a necessary condition is to reach beyond the critical Lundquist number, which is of the order of $10^{4}$ (Biskamp, 1986; Loureiro et al., 2007; Bhattacharjee et al., 2009; Gannouni et al., 2023). In our model, we find a large presence of plasmoids in the reconnecting current sheets when the refinement level is increased from 5 to 7 levels, i.e., when the local grid spacing is reduced from 1 Mm to 0.25 Mm.
V Synthetic In-Situ observation at streamers